I cant figure out this problem.
Im not sure what key is for absolute value so [] mean absolute value.
[2x-4]+7=15
please solve this soon.
Thank you

Answer by stanbon(48551)

[2x-4]+7=15
|2x-4| = 8
Since the result is 8, that 2x-4 must be 8 or -8
EQUATIONS
2x-4 = 8 or 2x-4 = -8
2x = 12 or 2x = -4
x = 6 or x = -2
====================
Cheers,
Stan H.